# UniHDA: Towards Universal Hybrid Domain Adaptation of Image Generators

### 23 Jan 2024 v1

@@ Abstract
Generative domain adaptation has achieved remarkable progress, enabling us to adapt a pre-trained generator to a new target domain. However, existing methods simply adapt the generator to a single target domain and are limited to a single modality, either text-driven or image-driven. Moreover, they are prone to overfitting domain-specific attributes, which inevitably compromises cross-domain consistency. In this paper, we propose UniHDA, a unified and versatile framework for generative hybrid domain adaptation with multi-modal references from multiple domains. We use CLIP encoder to project multi-modal references into a unified embedding space and then linear interpolate the direction vectors from multiple target domains to achieve hybrid domain adaptation. To ensure the cross-domain consistency, we propose a novel cross-domain spatial structure (CSS) loss that maintains detailed spatial structure information between source and target generator. Experiments show that the adapted generator can synthesise realistic images with various attribute compositions. Additionally, our framework is versatile to multiple generators, e.g., StyleGAN2 and Diffusion Models.
Figure 1: Application of UniHDA. Given a pre-trained source generator and multiple target domains (either in text or image modalities), UniHDA adapts the generator to a hybrid target domain that blends all characteristics and maintains strong consistency with source domain.
@@ 1 Introduction
Figure 2: Linear interpolation between multi-modal direction vectors. We represent the domain shift by the direction vector from source embedding to the target (e.g., Crying or Happy). Linear interpolation of them during training will result in a smooth traversal.
Benefiting from tremendous success of modern image generator [8, 1, 31, 25], generative domain adaptation has achieved remarkable progress during the past few years. Typically, it aims to adapt a pre-trained generator to a new target domain while preserving the variation in the source domain, e.g., from the human domain to the baby domain. Depending on the modality of references, generative domain adaptation can be categorized into two schools: text-driven [5, 17] and image-driven [15, 14, 19, 40, 35, 16].
Despite their promising results, there are still some limitations. A major limitation of existing methods is that they only support adaptation from a source domain to individual target domain. These methods fail to directly adapt a human generator to more practical real-world scenarios like angry person with the style of sketch or a combination of Johnson and Elena (Fig. 1). For more general purposes, previous methods also fail with multi-modal references, e.g., smiling Joker given the non-smiling Joker image and the smile text.
Another limitation of existing approaches is that, in the pursuit of faithful adaptation, they are prone to overfit domain-specific attributes, especially in image-driven scenarios [19]. This inevitably compromises cross-domain consistency and impedes the inheritance of the diversity from the source generator.
In this paper, we propose UniHDA, a Universal framework for generative Hybrid Domain Adaptation with multi-modal references from multiple domains. UniHDA facilitates the reference of individual image and text prompt simultaneously and blends the attributes from diverse target domains to create a hybrid domain. To enable multiple modalities, we leverage pre-trained CLIP [24] to project multi-modal references into a unified embedding space and represent the domain shift by the direction vector from the source embedding to the target embeddings.
To achieve hybrid domain adaptation, we draw inspiration from the compositional capabilities in the latent space of StyleGAN [6, 27, 36]. We demonstrate that a semantically meaningful linear interpolation between two direction vectors in CLIP’s embedding space can uncover favorable compositional capabilities, as shown in Fig. 2. In light of this intriguing finding, we linearly interpolate direction vectors of multiple target domains to obtain the direction vector corresponding to the hybrid domain that semantically integrates attributes from all target domains.
@@
We conduct experiments for a wide range of source and target domains to validate the effectiveness of our method. Results demonstrate that the adapted generator can synthesise realistic images with various attribute compositions. In addition, we show that UniHDA is agnostic to the type of generators (e.g., StyleGAN2 [8, 9, 10] and Diffusion models [7, 11]). Our contributions are as follows:
  • We propose a unified and versatile framework for generative domain adaptation, which enables multi-modal references and hybrid target domain, e.g., text-text, image-image, and image-text.
  • We demonstrate strong compositional capabilities of direction vectors in CLIP’s embedding space. Taking advantage of it, we propose to linearly interpolate the direction vectors for hybrid domain adaptation.
  • We propose a cross-domain spatial structure loss to maintain fine-grained spatial structure information between source and target generator, which significantly improves cross-domain consistency.
  • We demonstrate successful multi-modal hybrid domain adaptation to diverse new domains for various generators and illustrate its advantage over other methods.
@@ 2 Related Work
Text-driven Generative Domain Adaptation. Text-driven domain adaptation involves using a textual prompt to shift the domain of a pre-trained model toward a new domain. For example, Style-NADA [5] presents a local direction CLIP [24] loss to align the embeddings of the generated images and text. Based on Style-NADA, Domain Expansion (DE) [18] proposes to expand the generator to jointly model multiple domains with texts.
Image-driven Generative Domain Adaptation. Image-driven generative domain adaptation refers to the adaptation of a pre-trained image generator to a new target domain using a limited number of training images. Due to the scarcity of training images, prior researches  [15, 14, 19, 40, 35, 16, 33, 39, 41] often integrate additional regularization terms to prevent overfitting. For instance, CDC [19] introduces the instance distance consistency loss to maintain the distance between different instances in the source domain. Although these researches have made significant strides in generative domain adaptation, they heavily rely on the discriminator, making it challenging to handle hybrid domain adaptation.
Generative Hybrid Domain Adaptation. To combine multiple domains and achieve hybrid domain adaptation, several domain adaptation methods propose to train a separate generative model per domain and combine their effects in test-time, e.g., Style-NADA [5]. Although having potential for hybrid domain adaptation, it doubles the model size and training time due to the requirement for training multiple models separately. Domain Expansion (DE)  [18] proposes to expand the generator to jointly model multiple domains via decompose latent space. However, it requires the source dataset (e.g., FFHQ [8]) for regularization, which significantly increases training time. Furthermore, it relies on the semantic latent space of the generator (e.g., StyleGAN [8] and DiffAE [23]) for hybrid domain adaptation, limiting its applicability to a broader range of generators. Recently, FHDA [13] proposes few-shot hybrid domain adaptation and introduces a directional subspace loss to achieve it. Differently, we focus on multi-modal references with the text and 1-shot image, which offers greater flexibility and versatility.
@@ 3.1 Multi-Modal Hybrid Domain Adaptation
For more general purposes, we propose multi-modal hybrid domain adaptation. Given $N$ domains $\{\mathcal{T}_{i}\}_{i=1}^{N}$ with 1-shot image $\{Y_{i}\}$ and $M$ domains $\{\mathcal{T}_{j}\}_{j=1}^{M}$ with the text prompt $\{P_{j}\}$, it aims to adapt the source generator $G_{\mathcal{S}}$ to $G_{\mathcal{T}}$ that models the hybrid domain $\mathcal{T}=\{\mathcal{T}_{i}\}\cup\{\mathcal{T}_{j}\}$ and generates images with integrated characteristics. To the end, we introduce UniHDA, a unified and versatile framework for multi-modal hybrid domain adaptation, as shown in Fig. 3.
@@ 3.2 Multi-modal Direction Loss
To enable multiple modalities, we leverage pre-trained CLIP [24] to encode text-image references into a unified semantic embedding space. Drawing inspiration from CLIP-based methods [5, 38, 12, 11], we represent the domain shift as the direction vector $\Delta f_{dom}$ from the source embedding and the target embedding. For image reference $Y_{i}$ and its CLIP embedding $f_{i}$, the domain shift is calculated by
@@ 3.3 Linear Composition of Direction Vectors
In light of this intriguing finding, we employ linear interpolation on the direction vectors of multi-modal target domains, to derive a consolidated direction vector representing the hybrid domain that seamlessly integrates all attributes semantically. For given domain coefficients $\{\alpha_{i}\}$ and $\{\alpha_{j}\}$, we obtain the direction vector
@@ 3.4 Cross-domain Spatial Structure Loss
where $v_{i}^{t}$ and $v_{j}^{s}$ are the $i$-th and $j$-th tokens in the last layer of DINO-ViT from $G_{\mathcal{T}}$ and $G_{\mathcal{S}}$ respectively. The dot mark $\cdot$ represents cosine similarity.
@@
Table 1: Quantitative results for image-image domain adaptation.
Method Taylor-Elena Hulk-Wooden Johnson-Comic CS-I ($\uparrow$) SCS ($\uparrow$) CS-I ($\uparrow$) SCS ($\uparrow$) CS-I ($\uparrow$) SCS ($\uparrow$) FHDA 0.685 0.576 0.635 0.659 0.640 0.679 NADA 0.684 0.579 0.624 0.575 0.647 0.642 Ours 0.699 0.738 0.649 0.707 0.656 0.764
Method Baby-Red hair Baby-Happy Big-Blue eyes
CS-T ($\uparrow$) SCS ($\uparrow$) CS-T ($\uparrow$) SCS ($\uparrow$) CS-T ($\uparrow$) SCS ($\uparrow$)
DE 0.163 0.638 0.160 0.580 0.195 0.662
NADA 0.179 0.661 0.170 0.642 0.186 0.731
Ours 0.186 0.744 0.175 0.757 0.197 0.765
Table 1: Quantitative results for image-image domain adaptation.
Table 2: Quantitative results for text-text domain adaptation.
@@ 4.1 Experimental Setting
Methodology. We demonstrate the versatility of UniHDA on multi-modal hybrid domain adaptation, i.e., text-text, image-image, and image-text. To show the generator-agnostic nature of UniHDA, we apply it to two well-known generators, i.e., StyleGAN2 [9] and Diffusion model [11]). Following previous generative domain adaptation literatures  [5, 38, 15, 14, 18, 19, 40, 35, 16], we use StyleGAN2 for comparisons in most experiments.
Datasets. We conduct the experiments for a wide range of source and target domains to validate the effectiveness of UniHDA. Following previous work, we consider FFHQ [8], AFHQ Dog [3], and LSUN Church [37] as the source domains. We adapt the pre-trained models to diverse hybrid domains driven by the text prompt and 1-shot image, where we set the domain coefficients in Eq. 4 as 0.5. Unless stated otherwise, we operate on 1024 × 1024 resolution images for both source and target domains.
Evaluation Metrics. One important aspect to evaluate generative domain adaptation is the preservation of domain-specific characteristics. Following  [26], we use CLIP Score (CS-T and CS-I) for text-text and image-image adaptation respectively. Concretely, CS-T is measured by the average cosine similarity between prompt and generated images’ embedding. CS-I is the average pairwise cosine similarity between CLIP embeddings of generated and real images. Here we use average CS-T or CS-I of multiple domains. Another important evaluation is the cross-domain consistency of the source domain. To measure it, we adopt Structural Consistency Score (SCS) [35] to evaluate the spatial structural consistency between source and target generator.
@@ 4.2 Image-image Hybrid Domain Adaptation
Fig. 4 shows the qualitative results of baselines and UniHDA for image-image adaptation. As shown in the figure, FHDA [13] suffers from severe model collapse and generates images with a limited diversity due to the scarcity of image references. While NADA [5] mitigates overfitting to a certain extent, its cross-domain consistency remains poor, resulting in the generation of similar images. In contrast, UniHDA maintains strong consistency and effectively generates images with characteristics of the hybrid domain.
We also quantitatively compare UniHDA with baselines. As shown in Tab. 2, ours clearly outperforms other methods. We achieve better CS-I and SCS, indicating that generated images effectively integrate domain-specific attributes and preserve primary characteristics of the source domain.
@@ 4.3 Text-text Hybrid Domain Adaptation
Fig. 5 shows the qualitative results of the baselines and UniHDA for text-text adaptation, all of which start from the same source domain FFHQ [8] to the combinations of individual domains. Since the adaptation is conducted solely along one projection direction of the latent code, Domain Expansion (DE) [18], to some extent, does not fully capture the characteristics of the target domain, e.g., baby (Row 1 and Row 2). Furthermore, DE does not maintain robust cross-domain consistency, e.g., the chin of the person in the upper-right corner and background artifacts in the Row 3. This is primarily due to the characteristics bias introduced when projecting the latent code to the base subspace, making the images appear less authentic. The main problem of NADA [5] is overfitting. Hard-to-learn characteristics, e.g., baby (Row 1 and Row 2) and big eyes (Row 3) may be overshadowed by other overfitted ones. In contrast, UniHDA (Ours) exhibits desirable performance to generate images with integrated characteristics while maintaining robust consistency with the source domain.
Similar to Sec. 4.2, we also compare UniHDA with the baselines quantitatively. As shown in Tab. 2, ours clearly outperforms the baselines, which are consistent with qualitative results in Fig. 5. For CS-T, UniHDA significantly outperforms other methods, indicating that generated images effectively integrate multiple characteristics from distinct domains. Furthermore, UniHDA achieves better SCS, which effectively maintains cross-domain consistency compared with baseline methods.
Method Extra Data Dependency Model Dependency 2-domain 10-domain
@@
DE Source dataset $*{}^{*}$ 24M 20h 24M 20h
NADA - Multi-Models 48M 4min 240M 20min
Ours - Single-Model 24M 2min 24M 2min
  • *** indicates that DE needs a semantic latent space of the generator.
Table 3: Comparison with DE [18] and NADA [5]. For extra data dependency, DE needs source dataset (e.g., FFHQ [8]) that significantly increases the training time. Besides, DE requires a semantic latent space, which limits the generalization for more generators. For hybrid $N$-domains, NADA utilizes $N$ adapted models and necessitates $N$ times the model size and training time.
Figure 6: Image-text hybrid domain adaptation. Compared with NADA [5], UniHDA maintains strong cross-domain consistency.
Figure 7: Results of UniHDA for image-text hybrid domain adaptation on dogs [3], and churches [37]. The adapted generator can generate visually pleasant images with robust cross-domain consistency.
Figure 8: Results of UniHDA with DiffusionCLIP [11]. We replace the training objective of DiffusionCLIP with our proposed $\mathcal{L}_{direct}$ and $\mathcal{L}_{\text{CSS}}$. Results demonstrate that UniHDA is agnostic to the type of generators, allowing for its broader application on the diffusion model.
Figure 9: Ablation for our proposed $\mathcal{L}_{\text{CSS}}$ on single domains, which significantly alleviates overfitting and improves cross-domain consistency.

### 15 Mar 2024 v2

# UniHDA: A Unified and Versatile Framework for Multi-Modal Hybrid Domain Adaptation

@@ Abstract
Recently, generative domain adaptation has achieved remarkable progress, enabling us to adapt a pre-trained generator to a new target domain. However, existing methods simply adapt the generator to a single target domain and are limited to a single modality, either text-driven or image-driven. Moreover, they cannot maintain well consistency with the source domain, which impedes the inheritance of the diversity. In this paper, we propose UniHDA, a unified and versatile framework for generative hybrid domain adaptation with multi-modal references from multiple domains. We use CLIP encoder to project multi-modal references into a unified embedding space and then linearly interpolate the direction vectors from multiple target domains to achieve hybrid domain adaptation. To ensure consistency with the source domain, we propose a novel cross-domain spatial structure (CSS) loss that maintains detailed spatial structure information between source and target generator. Experiments show that the adapted generator can synthesise realistic images with various attribute compositions. Additionally, our framework is generator-agnostic and versatile to multiple generators, e.g., StyleGAN, EG3D, and Diffusion Models.
Keywords:
Generative Domain Adaptation Multi-Modal Adaptation Hybrid Domain Adaptation Generative Models
Figure 1: Given a pre-trained source generator and multiple target domains, UniHDA adapts the generator to a hybrid target domain that blends all characteristics at once and maintains robust cross-domain consistency. UniHDA supports both image and text modalities and is versatile to multiple generators.
@@ 1 Introduction
Figure 2: Existing methods like NADA [8] fail to maintain consistency with the source domain for hybrid domain adaptation, resulting in overfitting to the limited references and impeding the inheritance of the diversity in the source domain.
Figure 3: Linear interpolation between multi-modal direction vectors. We represent the domain shift by the direction vector from source embedding to the target (e.g., Crying or Happy). Linear interpolation of them during training will result in a smooth traversal. The coefficients for the right domain are respectively 0, 0.2, 0.4, 0.6, 0.8, and 1, while for the left domain, they are set inversely.
Benefiting from tremendous success of modern image generator [12, 3, 43, 35], generative domain adaptation has achieved remarkable progress during the past few years. Typically, it aims to adapt a pre-trained generator to a new target domain while preserving the variation in the source domain, e.g., from the human domain to the baby domain. Depending on the modality of references, generative domain adaptation can be categorized into two schools: text-driven [8, 27, 23, 40, 19, 14, 15, 2, 56, 24] and image-driven [25, 22, 29, 53, 47, 26, 45, 52, 54, 51, 55, 1, 7, 49, 17].
Despite their promising results, there are still some limitations. A major limitation is that they only support adaptation from a source domain to individual target domain. These methods fail to directly adapt a human generator to more practical real-world scenarios like person with red hair and blue eyes or a combination of Taylor and Elsa (Fig. 1). For more general purposes, previous methods also fail with multi-modal references, e.g., smiling Joker given the non-smiling Joker image and the smile text.
Another limitation of existing approaches is that, they are prone to overfit domain-specific attributes, especially in hybrid domain adaptation (Fig. 2). The main reason is that they do not maintain well consistency with the source domain, e.g., posture and identity. This results in overfitting to the limited references and losing the diversity in the source domain.
In this paper, we propose UniHDA, a Unified and versatile framework for generative Hybrid Domain Adaptation with multi-modal references from multiple domains. UniHDA facilitates the references of individual image and text prompt simultaneously and blends the attributes from target domains to create a hybrid domain. To enable multiple modalities, we leverage pre-trained CLIP [34] to project multi-modal references into a unified embedding space and represent the domain shift by the direction vector from the source embedding to the target embeddings.
To achieve hybrid domain adaptation, we draw inspiration from the compositional capabilities in the latent space of StyleGAN [9, 38, 48]. We demonstrate that a semantically meaningful linear interpolation between direction vectors in CLIP’s embedding space can uncover favorable compositional capabilities (Fig. 3). In light of this intriguing finding, we linearly interpolate direction vectors of multiple target domains to obtain the direction vector corresponding to the hybrid domain that semantically integrates attributes from all target domains.
@@
We conduct experiments for a wide range of source and target domains to validate the effectiveness of our method. Results demonstrate that the adapted generator can synthesise realistic images with various attribute compositions. In addition, we show that UniHDA is agnostic to the type of generators, e.g., StyleGAN [12, 13, 11], EG3D [4], and Diffusion models [10, 16]. Our contributions are as follows:
  • We propose a unified and versatile framework for generative domain adaptation, which enables multi-modal references and hybrid target domain, e.g., text-text, image-image, and image-text. We demonstrate successful adaptation to diverse domains for various generators and illustrate its advantage over other methods.
  • We demonstrate strong compositional capabilities of direction vectors in CLIP’s embedding space. Taking advantage of it, we propose to linearly interpolate the direction vectors for multi-modal hybrid domain adaptation.
  • We propose a cross-domain spatial structure loss to maintain consistency with source domain. It is conducted in generator-agnostic embedding space which is versatile for various generators, e.g., StyleGAN, EG3D, and Diffusion models. To our knowledge, it is the very first trial in generative domain adaptation.
@@ 2 Related Work
Text-driven Generative Domain Adaptation. Text-driven domain adaptation [8, 27, 23, 40, 19, 14, 15, 2, 56, 24] involves using a textual prompt to shift the domain of a pre-trained model toward a new domain. For example, Style-NADA [8] presents a local direction CLIP [34] loss to align the embeddings of the generated images and text. Based on Style-NADA, Domain Expansion (DE) [28] proposes to expand the generator to jointly model multiple domains with texts.
Image-driven Generative Domain Adaptation. Image-driven generative domain adaptation [25, 22, 29, 53, 47, 26, 45, 52, 54, 51, 55, 1, 7, 49, 17] refers to the adaptation of a pre-trained image generator to a new target domain using a limited number of training images. Due to the scarcity of training images, prior researches often integrate additional regularization terms to prevent overfitting. For instance, CDC [29] introduces the instance distance consistency loss to maintain the distance between different instances in the source domain. DiFa [51] utilizes GAN inversion [42] to align the latent codes which helps inherit diversity from the source generator. Although these works have made significant strides in generative domain adaptation, they heavily rely on the discriminator or generator, making it challenging to handle hybrid domain adaptation and extend to other generators.
Generative Hybrid Domain Adaptation. To achieve hybrid domain adaptation, several domain adaptation methods propose to train a separate generative model per domain and combine their effects in test-time, e.g., Style-NADA [8]. Although having potential for hybrid domain adaptation, it doubles the model size and training time due to the requirement for training multiple models separately. Domain Expansion (DE)  [28] proposes to expand the generator to jointly model multiple domains via decompose latent space. However, it requires the source dataset (e.g., FFHQ [12]) for regularization, which significantly increases training time. Furthermore, it relies on the semantic latent space of the generator (e.g., StyleGAN [12] and DiffAE [33]) for hybrid domain adaptation, limiting its applicability to a broader range of generators. Recently, FHDA [20] proposes few-shot hybrid domain adaptation and introduces a directional subspace loss to achieve it. Differently, we focus on multi-modal references with the text and one-shot image, which offers greater flexibility and versatility.
@@ 3.1 Multi-Modal Hybrid Domain Adaptation
For more general purposes, we propose multi-modal hybrid domain adaptation. Given $N$ domains $\{\mathcal{T}_{i}\}_{i=1}^{N}$ with one-shot image $\{Y_{i}\}$ and $M$ domains $\{\mathcal{T}_{j}\}_{j=1}^{M}$ with the text prompt $\{P_{j}\}$, it aims to adapt the source generator $G_{\mathcal{S}}$ to $G_{\mathcal{T}}$ that models the hybrid domain $\mathcal{T}=\{\mathcal{T}_{i}\}\cup\{\mathcal{T}_{j}\}$ and generates images with integrated characteristics. To the end, we introduce UniHDA, a unified and versatile framework for multi-modal hybrid domain adaptation (Fig. 4).
@@ 3.2 Multi-modal Direction Loss
To enable multiple modalities, we leverage pre-trained CLIP [34] to encode text-image references into a unified semantic embedding space. Drawing inspiration from CLIP-based methods [8, 51, 18, 16], we represent the domain shift as the direction vector $\Delta f_{dom}$ from the source embedding to the target embedding. For image reference $Y_{i}$ and its CLIP embedding $f_{i}$, the domain shift is calculated by
@@ 3.3 Linear Composition of Direction Vectors
In light of this intriguing finding, we employ linear interpolation on the direction vectors of multi-modal target domains, to derive a consolidated direction vector representing the hybrid domain that semantically integrates all attributes. For given domain coefficients $\{\alpha_{i}\}$ and $\{\alpha_{j}\}$, we obtain the direction vector
@@ 3.4 Cross-domain Spatial Structure Loss
where $v_{i}^{t}$ and $v_{j}^{s}$ are the $i$-th and $j$-th tokens in the last layer of Dino-ViT from $G_{\mathcal{T}}$ and $G_{\mathcal{S}}$ respectively. The dot mark $\cdot$ represents dot product.
@@ 4.1 Experimental Setting
Methodology. We demonstrate the versatility of UniHDA on multi-modal hybrid domain adaptation, i.e., image-image, text-text, and image-text. To show the generator-agnostic nature of UniHDA, we apply it to three well-known generators, i.e., StyleGAN2 [13], Diffusion model [16], and EG3D [4]. Following previous generative domain adaptation literatures  [8, 51, 25, 22, 28, 29, 53, 47, 26], we use StyleGAN2 for comparisons in most experiments.
Datasets. We conduct the experiments for a wide range of source and target domains to validate the effectiveness of UniHDA. Following previous work, we consider FFHQ [12], AFHQ-Dog [5], and LSUN-Church [50] as the source domains. The resolutions of images in these datasets are respectively 1024, 512, and 256. We adapt the pre-trained models to diverse hybrid domains driven by the text prompt and one-shot image. To demonstrate the effect of the hybrid domain, we set the domain coefficients in Eq. 4 as 0.5.
Method Taylor-Elena Hulk-Wooden Johnson-Comic Average
CS-I ($\uparrow$) SCS ($\uparrow$) CS-I ($\uparrow$) SCS ($\uparrow$) CS-I ($\uparrow$) SCS ($\uparrow$) CS-I ($\uparrow$) SCS ($\uparrow$)
FHDA 0.685 0.576 0.635 0.659 0.640 0.679 0.630 0.661
NADA 0.684 0.579 0.624 0.575 0.647 0.642 0.628 0.639
Ours 0.699 0.738 0.649 0.707 0.656 0.764 0.642 0.769
Table 1: Quantitative results for image-image domain adaptation. We present the quantitative results corresponding to each case in Fig. 5. To further demonstrate the robustness of our method, we average the results for 25 cases (shown in Appendix).
Evaluation Metrics. One important aspect to evaluate generative domain adaptation is the preservation of domain-specific characteristics. Following  [36], we use CLIP Score (CS-T and CS-I) for text-text and image-image adaptation respectively. Concretely, CS-T is measured by the average cosine similarity between prompt and generated images’ embedding. CS-I is the average pairwise cosine similarity between CLIP embeddings of generated and real images. Here we use average CS-T or CS-I of multiple domains. For image-text adaptation, we use the average of CS-T and CS-I as the metric (CS).
Another important evaluation is the cross-domain consistency of the source domain. To measure it, we adopt Structural Consistency Score (SCS) [47] to evaluate the spatial structural consistency between source and target generator.
@@ 4.2 Image-image Hybrid Domain Adaptation
Fig. 5 shows the qualitative results of baselines and UniHDA for image-image adaptation, starting from the same source domain FFHQ [12] to the combinations of individual domains. As shown in the figure, FHDA [20] suffers from severe model collapse and generates images with a limited diversity due to the scarcity of image references. While NADA [8] mitigates overfitting to a certain extent, its cross-domain consistency remains poor, resulting in the generation of similar images. In contrast, UniHDA maintains strong consistency and effectively generates images with characteristics of the hybrid domain.
We also quantitatively compare UniHDA with baselines. As shown in Tab. 1, ours clearly outperforms them. For CS-I, UniHDA significantly outperforms other methods, indicating that generated images effectively integrate multiple characteristics from distinct domains. Furthermore, UniHDA achieves better SCS, which effectively maintains cross-domain consistency compared with baselines.
Method Baby-Red hair Baby-Happy Big-Blue eyes Average
CS-T ($\uparrow$) SCS ($\uparrow$) CS-T ($\uparrow$) SCS ($\uparrow$) CS-T ($\uparrow$) SCS ($\uparrow$) CS-T ($\uparrow$) SCS ($\uparrow$)
DE 0.163 0.638 0.160 0.580 0.195 0.662 0.167 0.634
NADA 0.179 0.661 0.170 0.642 0.186 0.731 0.159 0.552
Ours 0.186 0.744 0.175 0.757 0.197 0.765 0.176 0.707
Table 2: Quantitative results for text-text domain adaptation. We present the quantitative results corresponding to each case in Fig. 6. Similar to Tab. 1, we average the quantitative results for 25 cases (shown in Appendix).
Figure 7: Image-text hybrid domain adaptation. We compared our method with previous [8, 55, 51], UniHDA well captures the attributes of hybrid target domain and maintains strong cross-domain consistency with source domain. *** indicates that MTG and DiFa support multi-modalities by interpolating model parameters with NADA.
@@ 4.3 Text-text Hybrid Domain Adaptation
Fig. 6 shows the qualitative results for text-text adaptation. Since the adaptation is conducted solely along one projection direction of the latent code, Domain Expansion (DE) [28], to some extent, does not fully capture the characteristics of the target domain, e.g., baby (Row 1 and Row 2). Furthermore, DE does not maintain robust consistency, e.g., the chin of the person in the upper-right corner and background artifacts in Row 3. This is primarily due to the characteristics bias introduced when projecting the latent code to the base subspace, making the images less authentic. The problem of NADA [8] is overfitting. Hard-to-learn characteristics, e.g., baby (Row 1 and Row 2) and big eyes (Row 3) may be overshadowed by other overfitted ones. In contrast, UniHDA (Ours) exhibits desirable performance to generate images with integrated characteristics while maintaining robust consistency with the source domain.
Similar to Sec. 4.2, we also compare UniHDA with the baselines quantitatively. As shown in Tab. 2, ours clearly outperforms the baselines, which are consistent with qualitative results in Fig. 6. We achieve better CS-I and SCS, indicating that generated images effectively integrate domain-specific attributes and preserve primary characteristics of the source domain.
Table 3: Quantitative results for image-text adaptation. We average the quantitative results for 81 cases for FFHQ, 16 cases for AFHQ-Dog, and 16 cases for LSUN-Church (shown in Appendix).
Method FFHQ Dog Church CS ($\uparrow$) SCS ($\uparrow$) CS ($\uparrow$) SCS ($\uparrow$) CS ($\uparrow$) SCS ($\uparrow$) NADA 0.563 0.586 0.424 0.533 0.414 0.629 MTG 0.536 0.529 0.403 0.526 0.403 0.684 DiFa 0.548 0.681 0.413 0.683 0.407 0.711 Ours 0.565 0.742 0.430 0.796 0.414 0.781
Method Fidel. Diver. Corr. vs. NADA (I-I) 85.2 90.6 76.0 vs. NADA (T-T) 81.4 84.2 80.8 vs. NADA (T-I) 84.6 85.8 78.6
Table 3: Quantitative results for image-text adaptation. We average the quantitative results for 81 cases for FFHQ, 16 cases for AFHQ-Dog, and 16 cases for LSUN-Church (shown in Appendix).
Table 4: User study for fidelity, diversity, and reference corespondence (image or text) in hybrid domain adaptation. The value (%) represents the percentage of users who favor the images generated by our method over NADA.
Method Modality Generator Dependency Model Amount 2-domain 10-domain
@@
NADA Multi - $N$ 48M 4min 240M 20min
$*{}^{*}$ Multi StyleGAN $N$ 48M 4min 240M 20min
$*{}^{*}$ Multi StyleGAN $N$ 48M 4min 240M 20min
${}^{\dagger}$ Text Latent space 1 24M 20h 24M 20h
FHDA Image - 1 24M 3min 24M 3min
Ours Multi - 1 24M 2min 24M 2min
[…]
Figure 9: Results of UniHDA with DiffusionCLIP [16], which demonstrate UniHDA is agnostic to the type of generator, allowing for broader application on diffusion models.
Figure 10: Ablation of our proposed $\mathcal{L}_{\text{CSS}}$ on hybrid domain adaptation, which significantly alleviates overfitting and improves cross-domain consistency.